OpenStudy (anonymous):

Fan and Medal What number should be added to the expression x^2 + 4x to change it into a perfect square trinomial? A. 2 B. 4 C. 8 D. 1

OpenStudy (anonymous):

@SandeepReddy

OpenStudy (anonymous):

perfect square trinomial ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we have the highest power of x as 2 so adding any number can only give us, at it's best. a perfect square binomial !

OpenStudy (anonymous):

i think the answer is 4 as x^2 + 4x + c = 0 we have \[(a + b )^2 = a^2 + 2ab + b^2\] so here a is x and we need to know b so that the number we should add becomes b^2 so 2ab = 4x 2*x*b = 4x b = 4x/2x = 2 so that led us to b = 2 so number we should is b^2 = > 2 squared = 4

OpenStudy (anonymous):

but that only makes it perfect square binomial not trinomial